# TANQ: An open domain dataset of table answered questions

### v1

@@ 4 Baselines & Evaluation
For all experiments, we use the PaLM-$2$ Unicorn model (Anil et al., 2023b), GPT4444 (Brown et al., 2020), [^5] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo and Gemini Pro (Anil et al., 2023a). We evaluate all baselines in a few-shot setting with three examples provided in the prompt (see top left prompt in Figure 3). The closed book setup evaluates LLMs’ capabilities in extracting relevant information acquired during training to answer TANQ questions.
@@
In the oracle setup, we provide models source attribution for every cell of the answer table in form of oracle documents (i.e. text, tables, and infoboxes). Hence, the prompt consists of the TANQ question and multiple evidence sentences, infobox entries and/or tables (see bottom left prompt in Figure 3). To provide further context, we include the Wikipedia page title, for text evidence also sub-/section titles, from where the evidence was extracted. Moreover, we add randomly selected oracle documents from other questions with similar attributes. This makes the oracle setting more challenging and requires models to filter the correct evidence from all provided ones first.
@@
We compare model performance against human performance based on $50$ answer tables generated by annotators. We provided the annotators the same input prompt as the models in the oracle setting (i.e. TANQ questions and evidence for each answer table cell).
@@
We evaluate answer tables using the RMS metrics introduced by (Liu et al., 2023). RMS views tables as unordered collections of mappings from row/colum headers to values. Hence, the metric is invariant to transpositions and column/row permutations. It allows small errors between tables keys/values of target and reference tables using the Normalized Levenstein Distance (Biten et al., 2019) with a threshold, which we set to $0.5$. The metrics returns both precision and recall scores.
@@ 5 Results & Discussion
In this section, we discuss the following research questions: $(1)$ Is TANQ a challenging dataset for state-of-the-art models? $(2)$ How well can models answer TANQ questions in a close book setting relying on knowledge acquired during training vs. extracted from provided context (oracle)? $(3)$ How well do tool-augmented models perform on TANQ? $(4)$ What are particular challenges related to question types, reasoning skills, complexity of questions, and the size of desired answer tables found in TANQ? $(5)$ What are common failure cases of evaluated models?
@@ 5.1 Question types
In Table 5, we compare the performance of all oracle, close, and open book baselines on TANQ. We find GPT4444 consistently outperforming other models with overall F1111 scores of $29.1$ and $19.6$ in oracle and closed book setting, respectively. While outperforming the other baselines, a large gap remains compared to our human baseline of $48.8$. While Gemini Pro (oracle) demonstrates higher precision compared to GPT4444 (oracle), there is still a meaningful gap for recall scores, indicating that Gemini successfully generates answer “subtables” containing only some of the requested rows. This aligns with our findings in Table 10, showing “missing entities” as the most prevalent failure type in Gemini (oracle) answers. In the oracle setup, most models perform best on simple questions, followed by composition and intersection. In the closed book setting, we observe higher performance on intersection questions than other types.
@@ 5.2 Reasoning skills
PaLM-$2$ 23.2 18.7 17.2 18.1 15.5 14.6 19.9 20.9 18.8
Gemini 33.0 13.2 11.2 33.1 12.6 11.0 25.7 16.0 12.8
GPT4 35.6 22.7 20.8 59.8 11.5 10.6 - - -
@@
Tool LM 20.1 17.5 15.3 15.5 15.1 12.5 14.8 22.1 16.8
Table 7: Baseline performance by number of skills used to augment TANQ questions.
We further augmented TANQ questions to generate more challenging questions requiring further skill sets in addition to search. Next, we discuss insights gained from baseline evaluation on each skill (see Table 6).
@@
PaLM-$2$ 20.2 19.0 15.4 20.5 17.0 15.4 15.8 12.3 12.3 20.5 17.7 18.5
Gemini 31.6 14.4 11.2 31.4 11.2 9.8 42.9 5.5 5.0 26.8 8.0 9.4
GPT4 32.7 32.9 27.9 29.2 15.9 15.1 43.8 5.3 6.5 25.5 0.2 0.3
@@
Tool LM 16.0 19.2 14.3 15.6 14.4 12.3 14.0 7.7 6.2 3.8 2.0 2.4
Table 8: Baseline performance by number of requested relations in TANQ questions. This results in columns of answer tables as demonstrated in Figure 1. PaLM-$2$ outperforms other baselines as the number of relations increases.
PaLM-$2$ (oracle/close/open).
For PaLM-$2$, we observe that the model struggles particularly with questions requiring numeracy, i.e. approximation of numerical values, quantity conversion, calculations based on datetime attributes. This issue persists for the tool augmented model. While having the Python tool available for questions requiring calculations, the agent model often ignores this tool for calculations. Moreover, the tool augmented counterpart demonstrates better performance on filtering tables with numerical and entity conditions ($17.1$ and $16.4$ respectively) but there is still a meaningful gap for filtering based on datetime conditions.
Gemini Pro (oracle/close).
Comparing to other baselines, we find Gemini Pro struggling particularly with the filtering skill for both oracle and close book evaluation. Filtering the answer table based on an entity condition is seemingly more challenging for the model than the other filtering categories. Moreover, we observe in Table 6 a substantial decrease of the F1111 score for the skill “time calc” (i.e. calculations based on datetime attributes) when comparing oracle and close book evaluation and a substantial gap to other baselines in the oracle setting.
GPT4444 (oracle/close).
Across all skills except approximation, the GPT4444 (oracle) baseline outperforms its counterparts. In the oracle setup, GPT4444 struggles however with numerical filtering, quantity conversion, and approximation of numerical values. On the other hand, it outperforms other baselines in entity filtering and time calculation. In the closed book setting the model demonstrates decreased performance compared to other closed book baselines for multiple skills, such as numerical filtering, date-to-year conversion, and time calculation.
@@
Table 7 shows baseline performance by the number of skills required to generate the correct answer table. Naturally, we observe across baselines a decrease of performance as the number of skills increases and questions become more complex. However, for Gemini and PaLM-$2$ this gap is smaller than for GPT4444 (both oracle and closed book).
5.3 Relations in TANQ questions
Table 8 demonstrates baselines’ performance across the number of relations in the given question. For example, in Figure 1 two additional relations, i.e. composer and release year, are requested for the movies specified in the question. While all models show a performance decrease with an increasing number of relations, this gap is particularly significant for the tool augmented model where the F1111 score decrease by almost twelve points comparing questions requiring one relations vs. ten. Moreover, we observe for all baselines (oracle/close/open book) little performance drops comparing questions with one relation to those with three. The performance gaps increase as the number of relations increases further, i.e. from three to five and from five to ten.
@@ 5.3 Relations in TANQ questions
PaLM-$2$ 16.8 17.6 11.6 17.8 16.6 13.3 22.6 18.0 17.3
Gemini 41.5 8.1 3.5 31.9 12.1 8.5 34.3 13.1 12.2
GPT4 19.4 25.6 20.0 27.9 23.5 19.4 42.3 20.4 19.7
@@
Tool LM 5.9 17.8 7.4 13.8 19.6 13.5 19.2 13.7 13.1
Table 9: Baseline performance by length of answer tables. Short tables have up to three rows, medium table up to six, and long table seven or more rows. For our best-performing baseline, GPT4444 (oracle), we observe a significant F1111 decrease with increasing table size.
@@ 5.4 Answer table length
Table 9 gives an overview of model performance across different answer table lengths: short answer tables (up to three rows), medium (up to six), and long tables (seven or more rows). A model seemingly struggling with longer tables is GPT4444 (oracle). The F1111 score decreases by $-4.4$ points comparing short to medium size tables and by further $-7.5$ points for long tables. It is apparent that the model struggles with extracting relevant information from the growing list of provided oracle documents as the table size increases. This is obviously not the case for close book evaluation as no oracle documents are given as part of the input.
@@ 5.5 Failures types
We further study common failures cases of oracle baselines. The aim is to understand challenges related to TANQ when the model has access to necessary information, in form of oracle documents, for answer table generation.
@@
Therefore, we manually evaluate a subset of model predictions to identify common failure cases which we organise in the following categories: 1.)1.)1 . ) Relation missing: At least one relation which corresponds to a column in the answer table is missing; 2.)2.)2 . ) No table header for the answer table is generated; 3.)3.)3 . ) Filtering (attribute, numeric, dates): Filtering conditions ignored or not applied correctly; 4.)4.)4 . ) Hallucinated relations: The answer table has columns which were not requested in the question; 5.)5.)5 . ) Hallucinations other other types of hallucination than additional relations or table format; 6.)6.)6 . ) Missing entity: entities, which correspond to rows in the table, are missing; 7.)7.)7 . ) Non-table Answers: Instead of generating a table as an answer, the model returns simple text or output in a different form than the table format provided in the prompt; 8.)8.)8 . ) Partial Answers: For a given entity and attribute only a partial answer is provided (e.g. only one university a person attended for relation “alma mater” although more exist); 9.)9.)9 . ) Wrong Answer (cell): Given answer (in a cell) is not correct.
Comparing open and oracle baselines, our observations show that the best performing model, w.r.t least failures, is GPT4444 (oracle). The only significant failure category we identify for GPT4444 are missing entities in generated answer tables. The most present issue we observe for Gemini Pro (oracle) are missing entities, missing relations, resulting in answer tables with a subset of columns, non-table outputs, and missing headers of generated tables. Similarly, multiple aspects pose a challenge for PaLM-$2$ (oracle), including hallucinations, missing entities/relations, and applying filtering conditions correctly. For the tool augmented model, one failure type is particularly prominent, namely missing entities that is followed by non-table outputs.
@@ 6 Conclusion
Appendix A TANQ Dataset
A.1 Prompt templates
Figure 4: Prompt used for evidence evaluation. We prompt a language model to evaluate the extracted evidence in a natural language inference setting. The LLM labels the input statements as “verifiable” or “not verifiable” based on evidence provided in form of a sentence, table or infobox.
Figure 5: Rephrasing questions increases the naturalness of template-based extension questions. To ensure the question meaning is preserved during rephrasing, we add structured annotations in parenthesis to questions with the name of each relation.
This section lists prompt templates we used for different steps of the TANQ dataset pipeline: evidence evaluation (Figure 4) and question rephrasing (Figure 5).

### v3

@@ 4 Baselines & Evaluation
[…]
@@
In the oracle setup, we provide models source attribution for every cell of the answer table in form of oracle documents (i.e. text, tables, and infoboxes). Hence, the prompt consists of the TANQ question and multiple evidence sentences, infobox entries and/or tables (see bottom left prompt in Figure 3). To provide further context, we include the Wikipedia page title, for text evidence also sub-/section titles, from where the evidence was extracted. Moreover, we add randomly selected oracle documents from other questions with similar attributes. This makes the oracle setting more challenging and requires models to filter the correct evidence from all provided ones first. We exclude models with a context length of less than $4k$ from the oracle evaluation. The evidence samples require a longer context to be fully displayed. Otherwise, performance may decrease due to the input being cut off.
@@
We compare model performance against human performance based on $100$ answer tables generated by annotators. We provided the annotators the same input prompt as the models in the oracle setting (i.e. TANQ questions and evidence for each answer table cell).
@@
To evaluate answer tables, we adopted a version of the relative mapping similarity (RMS) metrics introduced by Liu et al. (2023). RMS views tables as unordered collections of mappings from row/colum headers to values. Hence, the metric is invariant to transpositions and column/row permutations. It allows small errors between tables keys/values of target and reference tables using the Normalized Levenstein Distance (Biten et al., 2019). The metrics returns both precision and recall scores.
To evaluate the generated answer tables, we first converted table entries into a list of triplets. Each triplet consists of $(i)$ the entity name (given in the first column of the answer table), $(ii)$ the relation names (i.e., column names), and $(iii)$ the content of a table cell. If a table cell contains multiple values, such as in column Genre in Table 7, the cell content is split into multiple triplets with the same entity name and attribute name, but with different values extracted from the cell. For example, given Table 7, the resulting splitted triplets contains multiple triplets for genre: {(Evita, duration, 129 min), (Evita, genre, biographical film) (Evita, genre, musical), …}.
Film Duration Genre Publication
Evita 129 min biographical film, musical 1996
Table 7: Example TANQ answer table with multiple entries in a single cell, i.e., “biological film, musical”.
After generating triplet lists for the gold table and the model-generated answer table, we calculate the similarity between them using relative distance for numbers and Normalized Levenshtein distance for text across each field in the triplet. Each triplet in the target is matched with its closest triplet in the prediction greedily and thus we can compute weighted precision and recall, in the same manner as Liu et al. (2023). The evaluation code will be released.
@@ 5 Results & Discussion
In this section, we address key research questions: $(1)$ Is TANQ a challenging dataset for state-of-the-art models? $(2)$ How do models perform in a closed book setting compared to using external context (oracle)? $(3)$ How effective are tool-augmented models on TANQ? $(4)$ What challenges arise from different TANQ specifications, e.g., question types, reasoning skills, etc.? (5) What are the common failure cases of the evaluated models?
@@ 5.1 Question types
[…]
5.2 Evaluation with diverse prompts
In addition to the evaluations discussed in Section 5, we further assessed the baselines to examine the impact of $(i)$ instruction tuning and $(ii)$ demonstration selection on the models’ performance on TANQ. Table 15 compares the performance of baselines across different numbers of demonstrations in the input prompt (i.e., one-, three-, and five-shot), while Table 16 provides an overview of how different instruction styles affect the performance of Gemini Flash.
[…]
When comparing the performance of the top-performing oracle baseline, Gemini Flash, across different instruction styles, we find that the detailed (B), step-by-step (C), and simple (D) instruction styles yield similar performance (see Figure 6). In contrast, using an empty instruction (A) - where we provide only input examples without additional details - results naturally in decreased performance (see Table 16). For our evaluation in Section 5, we used the simple instruction style, as additional details as per instruction B and C did not contribute to observable performance improvements on the evaluated models.
@@ 5.3 Reasoning skills
PaLM-$2$ 52.8 46.2 47.4 52.0 46.9 47.9 57.0 47.1 50.5
Gemma 37.9 25.8 26.9 42.4 28.9 30.3 42.5 26.2 29.5
Gemini Pro 46.9 30.4 31.5 46.4 32.8 34.4 50.1 30.2 31.0
Gemini Flash 35.6 27.8 29.9 37.4 28.3 30.6 39.5 23.2 25.9
[…]
@@
Tool LM 40.7 33.5 35.0 46.4 35.6 39.8 - - -
Table 8: Baseline performance by number of skills required to successfully answer the TANQ questions.
Table 6 presents the performance of models on more challenging questions requiring further skills such as filtering and time conversion.
@@
PaLM-$2$ 53.7 47.4 48.4 51.8 45.2 46.0 54.2 47.7 48.8 62.4 47.7 48.9
Gemma 38.5 26.3 27.6 35.6 25.6 26.5 47.6 30.5 31.1 42.0 35.8 37.8
Gemini Pro 45.6 32.6 33.5 47.7 30.8 31.4 51.1 29.4 29.9 43.9 39.2 40.9
Gemini Flash 36.6 28.6 30.7 37.7 28.4 31.0 35.3 26.9 28.6 36.1 26.1 29.4
[…]
@@
Tool LM 31.6 32.0 30.2 42.4 33.7 36.7 34.0 32.9 33.0 18.4 17.8 18.1
Table 9: Baseline performance by number of relations in TANQ questions (i.e., corresponds to number of columns in the answer table). Most baselines demonstrate a stable performance as relations increase.
Gemini.
Overall, questions requiring filtering with numerical conditions poses the biggest challenge for Gemini models resulting in a performance drop compared to other skills. For example, Gemini Flash’s performance drops from $60.7$ F1111 (overall) to $55.3$ in oracle evaluation.
PaLM-$2$.
For PaLM-$2$, we observe that the model struggles particularly with questions requiring numeracy, i.e. approximation of numerical values, quantity conversion, calculations based on datetime attributes, and filtering with numerical conditions. This challenge does not persist in the open book baseline where PaLM-$2$ is augmented with a calculator tool. Despite these limitations, the PaLM-$2$ model outperforms all other baselines across all skills in the closed book setting.
[…]
For GPT-$4o$, no clear patterns are observed regarding its limitations in specific reasoning skills. In both the oracle and closed book scenarios, the model struggles with date-to-year conversions and the approximation of numerical values. However, it outperforms other closed book baselines in time calculations.
@@
[…]
@@
PaLM-$2$ 56.4 43.9 43.7 54.5 45.0 45.9 51.9 47.3 48.5
Gemma 35.8 25.8 25.6 37.6 25.1 25.9 39.4 27.0 28.4
Gemini Pro 47.8 29.6 29.7 45.9 31.5 32.7 47.0 30.5 31.6
Gemini Flash 32.9 25.7 25.6 36.3 27.9 29.4 37.1 28.9 31.4
[…]
@@
Tool LM 23.2 29.2 23.0 33.3 31.7 30.9 43.3 34.3 36.5
Table 10: Baseline performance by length of answer tables. Short tables have $<3$ rows, medium table up to six, and long table $>7$ rows. Most baselines show little performance variation across table sizes.
5.4 Relations in TANQ questions
[…]
@@ 5.5 Answer table length
Table 10 gives an overview of model performance across different answer table lengths: short answer tables (up to three rows), medium (up to six), and long tables (seven or more rows). Models that perform well on longer tables are GPT-$4o$, Gemini Flash, and PaLM-$2$. As table size increases, so does the number of rows the models can correctly retrieve, resulting in higher recall scores, while precision slightly drops for most oracle baselines models. It is apparent that the model struggles with extracting correct information from the growing list of provided oracle documents as the table size increases. This is obviously not the case for close book evaluation as no oracle documents are given as input.
@@ 5.6 Failures types
Hence, we further study common failures cases of oracle baselines. The aim is to understand challenges related to TANQ when the model has access to necessary information, in form of oracle documents.
@@
We manually evaluate a subset of model predictions to identify common failure cases, categorized as follows: 1.)1.)1 . ) Relation missing: At least one expected relation (column) is missing; 2.)2.)2 . ) No table header generated; 3.)3.)3 . ) Filtering: Conditions (attribute, numeric, dates) ignored or incorrectly applied; 4.)4.)4 . ) Hallucinated relations: Unrequested columns are added; 5.)5.)5 . ) Other hallucinations; 6.)6.)6 . ) Missing entity: Expected rows are missing; 7.)7.)7 . ) Non-table answers: Text or non-table format returned; 8.)8.)8 . ) Partial answers: Incomplete information for a given entity or attribute; 9.)9.)9 . ) Wrong answer (cell): Incorrect cell content.
[…]
5.7 Detailed evaluation of open book model
It is noteworthy that the open book, ReAct-based baseline performed worse than other baselines, also the closed book ones. To better understand the limitations of the open book baseline, we manually evaluated the reasoning chains generated by the model for $50$ TANQ samples. These chains outline the verbal reasoning processes and actions taken by the model in an interleaved manner. Our evaluation identified six different areas where the open book model struggles.
Retrieval of a complete entity list
The most prevalent issue, found in $19$ out of $50$ samples, was that the answer list did not contain all entities of the target table. Our review of the ReAct chain logs showed that the model often terminated the search for further entities after finding a list of entities. For example, it might start searching in the introduction section of a Wikipedia article and identify some movies directed by a particular director, but then fail to continue searching the rest of the page, missing additional films.
Generation of correctly-formatted answer tables.
We also found issues related to answer table formatting in seven out of fifty samples. For instance, in one case, the name column was incorrectly replicated into a table row (i.e., “Building” below) in Table 12.
name located in
building Old Main at the University of Arkansas Arkansas
Table 12: Example issue related to formatting answer tables: The first column is redundant.
Retrieval of attributes.
In six out of fifty samples, the model retrieved entities but failed to find attributes for all entities, resulting in partly empty rows (see Table 14).
Answering intersection questions.
A similar issue arises with intersection questions, which ask for entities that fulfill multiple conditions. For example, “What science fiction films were both written and directed by Steven Spielberg? What is their genre, publication date, and duration?” The open book model often produced tables that fulfilled only one of these conditions, such as listing films written by Steven Spielberg without considering if he also directed them. This problem occurred in five out of the fifty evaluated examples.
Answering composition questions.
As the name suggests, composition questions require the model to first retrieve a list of intermediate entities, which are then used to compile the final entity list needed for the answer table. For example, consider the question: “Who directed the film that Jules Feiffer wrote? What is their occupation, date of death, date of birth, country of citizenship, place of birth, and award received?” The model should first retrieve the films as intermediate entities to identify the requested directors. However, in four out of the fifty evaluated samples, the model generated answer tables that included intermediate entities only instead of the requested entities. For instance, the answer table for the above question looked as demonstrated in Table 13. The first row correctly shows the directors, but the second and third rows contain movies written by Jules Feiffer rather than their directors.
Applying filters to the generated table
Finally, the model commonly ignored filtering conditions specified in the input question. We observed this in ten out of fifty samples. For example, in the question: “Who was a mayor of Saint Paul, MN, and what was their occupation, position held, and date of birth? Filter the answer table for date of birth equal to or before 1829,” the resulting table included mayors born after $1829$, failing to apply the requested filter.
6 Future Work
Our results on TANQ highlight the need for qualitative benchmarks, metrics, and modeling approaches to address current limitations.
We initially used the RMS metrics from Liu et al. (2023) to evaluate table generation. However, RMS was developed for the chart-to-table conversion task, which involves mostly numerical tables. In contrast, TANQ tables contain significant text content, leading to variations such as expressing “USA” as “United States.” Another difference is the more diverse table structure, where cells often contain lists of values. We adapted the RMS metrics for TANQ (see Section 5). Our experiments indicate that this version of RMS strikes a right balance of providing signal, being simple to explain and implement, and fast to execute. However, further research on table evaluation is necessary to consider a broader range of formats.
Interestingly, we found that smaller models like Gemma ($9$B parameters) performed surprisingly well. While Gemma was not among the top models for any specific skill or question type in the closed book setting, it did not fall far behind models that are significantly larger, such as PaLM-$2$, Gemini Pro, and GPT-$4o$. This suggests that, despite its smaller size, models like Gemma and Gemini Flash offer promising results, though there is still room for improvement. While the size of Gemini Flash is undisclosed, it is known to be smaller than Gemini Pro, further highlighting that scale alone does not guarantee better performance on complex reasoning tasks.This raises an important question for future research: how can we further enhance the performance of smaller models?
Additionally, while tool-based LMs are popular, our results suggest that they may not be the optimal choice in domains where up-to-date knowledge is not required, especially for answering simple questions. For more complex questions involving intersecting information, tool-augmented LMs perform comparably to other baselines, but for simpler questions, they lag behind both oracle and closed book baselines.
Numeracy remains a challenge for state-of-the-art models. Despite advancements in numerical reasoning with LLMs, many models still struggle with questions requiring simple numerical skills such as filtering the content of a based on a numerical condition. While progress has been made in recent years (Imani et al., 2023; Akhtar et al., 2023; Chen et al., 2022), challenges remain.
Another challenge is that all evaluated models tend to generate incomplete answer tables, often missing some entities (rows) and relations (columns) from the target table. Future work should focus on developing methods to address these limitations and generate answer tables that capture the entire requested information. Addressing these issues and improving table evaluation metrics are important areas for future research.
@@ 7 Conclusion
Appendix A Modeling details
[…]